OpenStudy (anonymous):

Create a unique example of dividing a polynomial by a monomial and provide the simplified form. Explain, in complete sentences, the two ways used to simplify this expression and how you would check your quotient for accuracy.

OpenStudy (anonymous):

@ganeshie8 @phi @jdoe0001 @whpalmer4

OpenStudy (anonymous):

can you help

OpenStudy (phi):

can you write down a monomial ?

OpenStudy (anonymous):

x

OpenStudy (phi):

can you write down a polynomial ?

OpenStudy (anonymous):

5x2+2x3?

OpenStudy (phi):

ok, but you should use ^ to show exponents. 5x^2 + 2x^3 normally we write the terms in order of biggest exponent to smallest, so this way \[ 2x^3 + 5x^2\] I was going to suggest multiplying the monomial by the polynomial, but your polynomial can be divided by \(x^2\) so try \[ \frac{ 2x^3 + 5x^2}{x^2} \]

OpenStudy (anonymous):

2x^5+5x^4? @phi

OpenStudy (phi):

you multiplied \[ x^2(2x^3 +5x^2)\] that can be "distributed" by multiplying each term inside the parens \[ x^2 \cdot 2x^3+ x^2 \cdot 5x^2\] when you multiply, you can change the order (for example, with numbers 2*3 is the same as 3*2... it works the same for letters) \[ 2\cdot x^2 \cdot x^3+ 5 \cdot x^2 \cdot x^2\] x^2 is short for x*x x^3 is short for x*x*x x^2 * x^2 is short for x*x * x*x*x or, using the short way, x^5 but you do not want to multiply you want to divide.

OpenStudy (phi):

**typo** x^2 is short for x*x x^3 is short for x*x*x x^2 * x^3 is short for x*x * x*x*x or, using the short way, x^5

OpenStudy (phi):

dividing is almost the same as multiplying \[ \frac{ 2x^3 + 5x^2}{x^2} \] divide x^2 into each term \[ \frac{2x^3}{x^2} + \frac{5x^2}{x^2} \]

OpenStudy (phi):

if you write x^3 the long way, as x*x*x and x^2 as x*x, the problem is \[ \frac{2 \cdot x \cdot x \cdot x }{ x \cdot x } + \frac{5 \cdot x \cdot x }{ x \cdot x }\]

OpenStudy (phi):

when you divide anything by itself you get 1 in other words, every time you can pair up an x in the top with an x in the bottom, they divide into themselves and give you 1

OpenStudy (phi):

not clear ?